# Input Reduction Enhanced LLM-based Program Repair

### v1

@@ 5.3. RQ2: Effectiveness of ReduceFix
C 56.5 76.3 80.0 56.7 ($+0.2\%$) 75.5 ($-0.9\%$) 78.3 ($-1.7\%$) 56.3 ($-0.2\%$) 77.6 ($+1.2\%$) 81.7 ($+1.7\%$)
D 44.5 59.1 62.5 48.9 ($+4.4\%$) 59.2 ($+0.1\%$) 60.0 ($-2.5\%$) 48.1 ($+3.6\%$) 61.0 ($+1.9\%$) 63.7 ($+1.2\%$)
E&F 35.0 53.2 58.3 32.0 ($-3.0\%$) 50.0 ($-3.1\%$) 51.7 ($-6.7\%$) 32.5 ($-2.5\%$) 51.4 ($-1.8\%$) 56.7 ($-1.7\%$)
Overall 45.2 62.5 66.5 46.1 ($+0.9\%$) 61.3 ($-1.2\%$) 63.0 ($-3.5\%$) 45.9 ($+0.6\%$) 63.1 ($+0.6\%$) 67.0 ($+0.5\%$)
[Experimental Results]: Across all 4 evaluated LLMs, providing the reduced failure‑inducing input consistently raises repair accuracy. Table 7 shows that the overall pass10 climbs from 20% to 25.5% for Qwen2.5‑Coder‑7B‑instruct, from 8.5% to 10.0% for GLM4‑9B‑chat, from 59.0% to 61.0% for Qwen‑Plus, and from 66.5% to 67.0% for DeepSeek‑V3 when the Reduced Test prompt replaces the Baseline prompt without any other change.
The benefit intensifies on more complex bugs. On D‑difficulty tasks, the Reduced Test prompt lifts Qwen2.5’s pass@10 by 44.8% and pushes GLM4‑9B‑chat up by 44.9%. Tasks with difficulties E&F see Qwen‑Plus rise from 13.9%, and DeepSeek‑V3 recover part of the loss incurred by the origin test input, moving from 51.7% to 56.7%.
In contrast, prompting the unreduced test case often hampers performance. GLM4‑9B‑chat’s overall pass@10 falls from 8.5% with no test case to 6.5% when the complete input is added, while DeepSeek‑V3 drops from 66.5% to 63.0% under the same switch. These observations confirm that a focused counterexample, as provided by ReduceFix, supplies the LLM with sufficient evidence to pinpoint the defect.
To test whether the improvement is due to randomness, we leverage MWW tests (Mann and Whitney, 1947; Wilcoxon, 1992) to further confirm the improvement, returning a two‑sided p‑value with $<0.05$; therefore, the null hypothesis of equal success probabilities between ReduceFix and the “Origin Test” is rejected, establishing that the gain is statistically significant.
[RQ‑2] [Findings]: (1) Supplying the full failing test often hurts repair accuracy; in multiple LLMs and difficulty levels, it performs worse than the no‑test Baseline. (2) The reduced test produced by ReduceFix is consistently better than Origin Test across all LLMs and difficulty groups, and it also outperforms the Baseline in every overall comparison. [Insights]: Trimming a long test while preserving its failure signal boosts APR performance, suggesting that controlled input compression may benefit other long‑context tasks that must balance prompt length with information retention.
@@ 5.4. RQ-3: Influence of Prompt Composition on Repair Performance
[Objective]: We investigate the distinct influence of two factors within ReduceFix: (i) length reduction (fewer tokens to keep the bug‐relevant text within the model’s attention span) and (ii) information selection (retaining the minimal concrete evidence that still exposes the defect).
[Experimental Design]: All settings that are unrelated to the prompt remain identical to RQ‑2: the benchmark, the Qwen 2.5 Coder‑7B‑instruct model, the decoding temperature of 0.8, and the sampling times of ten candidate patches per bug. Five prompt variants cover every combination of the two dimensions under study, and an additional control is included with no test information. The Baseline prompt contains only the problem statement and the buggy code. Diff Lines stays the same length but appends up to ten mismatched output lines, providing sparse evidence without increasing size. Reduced Test includes the full input–output pair of the minimized counterexample; it represents the joint action of length control and full information and is the default in ReduceFix. Origin Test swaps the reduced input for the unreduced failure‑inducing case, inflating the prompt to about 30.6 KB while holding informational content constant, thereby isolating the cost of extra length. Reduced + Origin concatenates the reduced and full tests, introduces redundant information as well as maximum length.
1### Problem Description
2{full problem text}
3### Your Incorrect Code
4“‘cpp
5{buggy code here}
6““
7### Error Summary (diff only)
8Line 1: Got ’42’, Expected ’43’
9Line 2: Got ’…’, Expected ’…’
10
11### Your Task
12Fix the code so that the diff disappears on all tests.
13Return only the complete corrected C++ program in a “‘cpp block.
Listing 4: Template for Diff Lines
For every bug, the LLM is called exactly once with each variant. After generation, we compile and run the candidate patches against the full hidden test suite and record pass@k for $k\in{1,5,10}$. Mean prompt length is reported alongside pass@k for each difficulty group (C, D, E&F).
[Experimental Results]: Table 8 confirms that the five prompt variants span more than an order of magnitude in length, from roughly 3 KB for Baseline and Diff Lines to 36 KB for Reduced + Origin. Repair accuracy tracks these length and content differences in a highly systematic way (Table 9). Across the full 200‑bug benchmark, Reduced Test delivers the best outcomes, reaching 25.5% pass@10. This score is 5.5 percentage points higher than the Baseline that omits test evidence and 5.5 points higher than the Origin Test that embeds the unreduced input.
@@
Strategy Mean (KB)
Baseline 3.1
Origin Test 30.6
Diff Lines 3.1
Reduced Test 6.4
Reduced and Origin Test 36.4
@@
Two clear patterns emerge. First, inserting the entire failure-inducing input without reduction inflates the prompt by roughly 30 KB and consistently depresses success, confirming that excessive length dilutes attention. Second, supplying only a terse diff helps little once problems become non‑trivial because the LLM lacks a concrete input–output correspondence. The conjunction of compact length and complete counter‑example information is therefore essential; either ingredient alone is insufficient. These findings demonstrate that ReduceFix’s prompt strategy offers the best trade‑off between information richness and context length, a conclusion likely to generalize to other long‑context code‑related tasks.
@@
5.5. RQ‑4: extend ReduceFix to other APR pipelines
[Objective]: We validate the extensibility of ReduceFix by adding it as a plug‑in to the state-of-the-art APR system ChatRepair and observing whether the straightforward replacement of the full failure‑inducing input with the reduced counter‑example leads to a higher patch accuracy.
[Experimental Design]: ChatRepair (Xia and Zhang, 2024) is a conversational repair framework that alternates between a user proxy and an LLM. The first user turn delivers the task description, the buggy C++ program, and a single failing test case. The assistant then proposes a patch, which the testing harness compiles and executes; its verdict (pass or fail) becomes the next user message. A run stops when a patch passes all tests or when the retry budget is exhausted. The original implementation exposes two key hyperparameters: MAX_RETRY, which limits the number of feedback rounds, and length, which decides how many previous turns are retained in the next prompt.
We evaluate ChatRepair (Xia and Zhang, 2024) on the built LFTBench benchmark under two prompting modes:
@@ 5.5. RQ‑4: extend ReduceFix to other APR pipelines
To isolate the effect of the initial prompt, we fix MAX_RETRY$=1$ (one feedback round) and set the conversation window size to length$=2$ turns. The LLM samples up to ten candidate patches per bug; we record pass@k for $k\!\in\!\{1,5,10\}$ and results are grouped by difficulty.
@@
[Experimental Results]: Table 10 indicates that replacing the original failure‑inducing input with the ReduceFix counter‑example improves ChatRepair across every difficulty level. The overall pass@10 rises from 30.5% to 37.0%, an absolute gain of 6.5 percentage points and a relative gain of 21.3%. Looking by difficulty, C-difficulty tasks improve from 41.7% to 45.0% (7.9% relative improvement), D-difficulty tasks from 37.5% to 46.2% (23.2% relative improvement), and the hardest E&F tasks from 10.0% to 16.7% (67.0% relative improvement). Similar lifts appear at pass@5 and pass@1, confirming that the benefit is robust across different sampling $k$. These results demonstrate that reducing the supplied failure-inducing test yields a measurable improvement without requiring any other adjustments to existing APR systems, such as ChatRepair.
@@
[RQ-4] [Findings]: Integrating ReduceFix with ChatRepair increases the overall pass@10 by 21.3% relative to the original configuration. [Insights]: ReduceFix can be adopted as a lightweight add‑on for any other APR system that leverages test cases, providing an immediate boost in repair effectiveness while requiring no changes to the existing APR pipeline.

### v2

@@ 5.3. RQ2: Effectiveness of ReduceFix
C 56.5 76.3 80.0 56.7 75.5 78.3 56.3 77.6 81.7
D 44.5 59.1 62.5 48.9 59.2 60.0 48.1 61.0 63.7
E&F 35.0 53.2 58.3 32.0 50.0 51.7 32.5 51.4 56.7
Overall 45.2 62.5 66.5 46.1 61.3 63.0 45.9 63.1 67.0
[Experimental Results]: Across all 4 evaluated LLMs, providing the reduced failure-inducing input consistently raises repair accuracy. Table 6 shows that the overall pass@10 climbs from 20% to 25.5% for Qwen2.5-Coder-7B-instruct, from 8.5% to 10.0% for GLM-4-9B-chat, from 59.0% to 61.0% for Qwen2.5-Plus, and from 66.5% to 67.0% for DeepSeek-V3 when the Reduced Test prompt replaces the Baseline prompt without any other change.
Gains are strongest on D-difficulty tasks. On D-difficulty tasks, the Reduced Test prompt lifts Qwen2.5’s pass@10 by 44.8% and pushes GLM-4-9B-chat up by 44.9%. Tasks with difficulties E&F see Qwen2.5-Plus rise from 51.7% to 55.0%, and DeepSeek-V3 recover part of the loss incurred by the origin test input, moving from 51.7% to 56.7%.
In contrast, prompting the unreduced test case often hampers performance. GLM-4-9B-chat’s overall pass@10 falls from 8.5% with no test case to 6.5% when the complete input is added, while DeepSeek-V3 drops from 66.5% to 63.0% under the same switch. These observations confirm that a focused counterexample, as provided by ReduceFix, supplies the LLM with sufficient evidence to pinpoint the defect.
To test whether the improvement is due to randomness, we leverage MWW tests (mann1947test; wilcoxon1992individual) to further confirm the improvement based on 4 LLMs, returning a two-sided p-value with <0.05<0.05; therefore, the null hypothesis of equal success probabilities between ReduceFix and the “Origin Test” is rejected, establishing that the gain is statistically significant. On LFTBench with Qwen2.5-Coder-7B-instruct, leveraging reduced test produced by ddmin-only attains 5.7% pass@1, 16.1% pass@5, and 22.0% pass@10, compared to 6.3%, 17.9%, and 25.5% for ReduceFix; the gap at pass@10 is also statistically significant under a two-sided MWW test (p ¡ 0.05).
On LFTBench-Py with Qwen2.5-Plus, Reduced Test lifts overall pass@10 to 75.0% versus 70.0% for Baseline and 60.0% for Origin, with the strongest gains on Medium tasks (Table 7).
Table 7. ReduceFix with Qwen2.5-Plus on LFTBench-Py by difficulty.
Difficulty Baseline (No Test) Origin Test Reduced Test
@1 @5 @10 @1 @5 @10 @1 @5 @10
C 34.8 62.2 66.7 38.9 62.2 66.7 38.3 64.5 66.7
D 53.9 69.1 75.0 52.5 62.2 62.5 61.3 84.4 87.5
E&F 35.8 66.7 66.7 18.3 40.3 50.0 35.2 61.5 66.7
Overall 42.7 66.3 70.0 38.2 55.6 60.0 46.5 71.5 75.0
[RQ-2] [Findings]: (1) Supplying the full failing test often hurts repair accuracy; in multiple LLMs and difficulty levels, it performs worse than the no-test Baseline. (2) ReduceFix is consistently better than Origin Test across all LLMs and difficulty groups, and it also outperforms the Baseline in every overall comparison. (3) When evaluating on LFTBench-Py, the improvement of ReduceFix is significant, indicating its cross-language robustness. [Insights]: Trimming a long test while preserving its failure signal boosts APR performance, suggesting that controlled input compression may benefit other long-context tasks that must balance prompt length with information retention.
@@ 5.4. RQ-3: Influence of Prompt Composition on Repair Performance
[Objective]: We investigate the distinct influence of two factors within ReduceFix: (i) length reduction (fewer tokens to keep the bug-relevant text within the model’s attention span) and (ii) information selection (retaining the minimal concrete evidence that still exposes the defect).
[Experimental Design]: All settings that are unrelated to the prompt remain identical to RQ-2: the benchmark, the Qwen 2.5 Coder-7B-instruct model, the decoding temperature of 0.8, and the sampling times of ten candidate patches per bug. Five prompt variants cover every combination of the two dimensions under study, and an additional control is included with no test information. The Baseline prompt contains only the problem statement and the buggy code. Diff Lines stays the same length but appends up to ten mismatched output lines, providing sparse evidence without increasing size. Reduced Test includes the full input–output pair of the minimized counterexample; it represents the joint action of length control and full information and is the default in ReduceFix. Origin Test swaps the reduced input for the unreduced failure-inducing case, inflating the prompt to about 30.6 KB while holding informational content constant, thereby isolating the cost of extra length. Reduced + Origin concatenates the reduced and full tests, introducing redundant information as well as maximum length.
For every bug, the LLM is called exactly once with each variant. After generation, we compile and run the candidate patches against the full hidden test suite and record pass@k for $k\in\{1,5,10\}$. Mean prompt length is reported alongside pass@k for each difficulty group (C, D, E&F).
@@
Strategy Mean Lines Mean KBs
Baseline 130.1 3.2
Origin Test 2381.4 30.5
Diff Lines 132.6 3.2
Reduced Test 514.0 6.6
Reduced and Origin Test 2637.7 36.4
[Experimental Results]: Table 8 confirms that the five prompt variants span more than an order of magnitude in length, from roughly 3 KB for Baseline and Diff Lines to 36 KB for Reduced + Origin. Repair accuracy tracks these length and content differences in a highly systematic way (Table 9). Across the full 200-bug benchmark, Reduced Test delivers the best outcomes, reaching 25.5% pass@10. This score is 5.5 percentage points higher than the Baseline that omits test evidence and 5.5 points higher than the Origin Test that embeds the unreduced input.
@@
Two clear patterns emerge. First, inserting the entire failure-inducing input without reduction inflates the prompt by roughly 30 KB and consistently depresses success, confirming that excessive length dilutes attention. Second, supplying only a terse diff helps little once problems become non-trivial because the LLM lacks a concrete input–output correspondence. The conjunction of compact length and complete counter-example information is therefore essential; either ingredient alone is insufficient. These findings demonstrate that ReduceFix’s prompt strategy offers the best trade-off between information richness and context length, a conclusion likely to generalize to other long-context code-related tasks.
@@
5.5. RQ-4: Extending ReduceFix to Other APR Pipelines
[Objective]: We validate the extensibility of ReduceFix by adding it as a plug-in to the state-of-the-art APR system ChatRepair and CREF, and observing whether the straightforward replacement of the full failure-inducing input with the reduced counter-example leads to a higher patch accuracy.
[Experimental Design]: We select ChatRepair (xia2024automated) and CREF (yang2024cref) for their open harnesses and drop-in reducer interfaces, which let us replace only the failing test while keeping all other logic unchanged. We fix the sampling budget (k) across settings to isolate the effect of reduced vs. unreduced tests. We keep $k=10$ across settings to equalize token budgets; increasing k raises absolute pass@k but does not affect our budget-controlled comparisons.
ChatRepair (xia2024automated) is a conversational repair framework that alternates between a user proxy and an LLM. The first user turn delivers the task description, the buggy C++ program, and a single failing test case. The assistant then proposes a patch, which the testing harness compiles and executes; its verdict (pass or fail) becomes the next user message. A run stops when a patch passes all tests or when the retry budget is exhausted. The original implementation exposes two key hyperparameters: MAX_RETRY, which limits the number of feedback rounds, and length, which decides how many previous turns are retained in the next prompt.
We evaluate ChatRepair and CREF on the built LFTBench benchmark under two prompting modes:
@@ 5.5. RQ-4: Extending ReduceFix to Other APR Pipelines
To isolate the effect of the initial prompt, we fix MAX_RETRY=1=1 (one feedback round) and set the conversation window size to length=2=2 turns. The LLM samples up to ten candidate patches per bug; we record pass@k for $k\!\in\!\{1,5,10\}$ and results are grouped by difficulty.
CREF (yang2024cref) is a conversational repair framework for programming tutors that iteratively refines a patch using tutor guidance and test feedback across turns. Since CREF expects a tutor-provided hint in its first turn but LFTBench does not provide such hints, we omit that turn and run a two-turn variant: we first supply the official AtCoder editorial as the high-level solution description, then provide the failure-inducing test case returned by the harness after validating the first-turn patch as the second-turn feedback.
@@
Table 11. Repair success of CREF with full versus reduced failure-inducing inputs on LFTBench.
Difficulty CREF CREF + ReduceFix
pass@1 pass@5 pass@10 pass@1 pass@5 pass@10
C 15.8 37.3 45.0 17.3 39.5 46.7
D 18.8 39.4 47.5 19.4 39.7 47.5
E&F 5.5 15.6 21.7 5.7 16.4 23.3
Overall 13.9 31.6 39.0 14.7 32.7 40.0
@@
[Experimental Results]: Table 10 indicates that replacing the original failure-inducing input with the ReduceFix counter-example improves ChatRepair across every difficulty level. The overall pass@10 rises from 30.5% to 37.0%, an absolute gain of 6.5 percentage points and a relative gain of 21.3%. Looking by difficulty, C-difficulty tasks improve from 41.7% to 45.0% (7.9% relative improvement), D-difficulty tasks from 37.5% to 46.2% (23.2% relative improvement), and the hardest E&F tasks from 10.0% to 16.7% (67.0% relative improvement). Similar lifts appear at pass@5 and pass@1, confirming that the benefit is robust across different sampling $k$. Likewise, CREF benefits from ReduceFix: overall pass@10 rises from 39.0% to 40.0%, with pass@5 from 31.6% to 32.7% and pass@1 from 13.9% to 14.7% on LFTBench (Table 11). These results demonstrate that reducing the supplied failure-inducing test yields a measurable improvement without requiring any other adjustments to existing APR systems.
@@
[RQ-4] [Findings]: Integrating ReduceFix with ChatRepair increases the overall pass@10 by 21.3% relative and with CREF by 2.6% to the original configuration. [Insights]: ReduceFix can be adopted as a lightweight add-on for any other APR system that leverages test cases, providing an immediate boost in repair effectiveness while requiring no changes to the existing APR pipeline.
5.6. RQ-5: Experiments on OSS-Fuzz
[Objective]: We aim to evaluate the effectiveness of ReduceFix in realistic software-level repair scenarios beyond C++ programming-exercise benchmarks. Specifically, we test whether ReduceFix can serve as a general, language-agnostic enhancement to LLM-based automated program repair by validating it on real OSS-Fuzz crash instances, assessing both the robustness of the proposed input-reduction mechanism in large, noisy industrial environments and its cross-language applicability across C, C++, and Bash buggy codes.
[Experimental Design]: We evaluate a guided repair approach for automated fixing of instances in OSS-Fuzz projects (googleAnnouncingOSSFuzzContinuous). We reuse the data and scripts prepared by ARVO (mei2024arvo) to extract ground-truth patches, failure-inducing inputs, codebases, and issue descriptions, and to evaluate generated patches using pre-built Docker images. The repair pipeline operates iteratively according to the settings in Table 2, except that the reduction phase uses a 600-second time limit to account for Docker test execution overhead. Each LLM invocation receives structured contextual information, including crash diagnostics, complete source files extracted from vulnerable repositories, and failure-inducing input in hexadecimal format. The prompt enforces strict output formatting, including SEARCH/REPLACE blocks with file paths and line numbers for automated patch generation, consistent with prior repository-level repair frameworks (yang2025enhancing; xia2024agentless). Each generated patch is immediately validated in the original Docker environment provided by ARVO. For ddmin-only, we use whitespace/newline tokens for line-oriented textual inputs as in Section 5.2; for binary inputs, we adopt fixed 256-byte blocks with a byte-level fallback. We report the best 1-minimal candidate found within the time budget.
To ensure comparability and avoid post-hoc cherry picking, we first filter instances from ARVO whose ground-truth patch modifies exactly one file and whose failure-inducing input exceeds 1 KB. From the pool that satisfies these constraints across five projects (FFmpeg, ImageMagick, Poppler, php-src, MuPDF), we then select the 12 instances with the smallest ground-truth patch size measured by unified diff lines. We compare the effectiveness of reduction across three strategies (ReduceFix, ddmin-only, and LLM-generated reduction) by reporting the reduction success rate and compression ratio. Finally, we evaluate the repair accuracy of three strategies (ReduceFix, Baseline without test, and Origin Test) using pass@1, pass@5, and pass@10.
[Experimental Results]: As shown in Table 12 and Table 13, on 12 OSS-Fuzz crashes spanning FFmpeg, Poppler, PHP-src, and MuPDF, ReduceFix preserved the crash during reduction in 91.7% of cases versus 75.0% for ddmin-only, a +22.3% relative gain and an absolute +16.7 points. Mean compression rose from 51.8% to 56.4%, a +8.9% relative gain. Meanwhile, ddmin-only failed outright on FFmpeg, while ReduceFix both preserved the crash and reduced the input by 35.4%. On PHP-src, compression improved from 85.6% to 97.7% (+14.1% relative) with crash preservation at 100% for both. On Poppler, preservation increased from 50.0% to 75.0% (+50.0% relative) and compression from 22.8% to 25.4% (+11.4% relative). On MuPDF, both preserved the crash, and compression was comparable (-0.8% relative for ReduceFix). The pure-LLM approach fails to reduce any of the failure-inducing test inputs, mainly because these inputs are long (exceeding 1 KB) and structurally complex (e.g., binary file), which LLMs cannot effectively simplify.
Under Docker-grounded validation with Qwen2.5-Plus, reduced tests improved end-to-end repair accuracy. On the 12-sample micro average, pass@1 rose from 17.5% to 19.2% (+9.7% relative), pass@5 from 25.0% to 29.2% (+16.8% relative), and pass@10 from 25.0% to 41.7% (+66.8% relative). Using the unreduced Origin Test as the baseline, reduced tests reached 19.2%, 29.2%, and 41.7%, which correspond to +108.7%, +97.3%, and +149.7% relative gains. By project, FFmpeg reached 100.0% pass@10 from 0%; MuPDF doubled pass@10 from 20.0% to 40.0%; PHP-src raised pass@1 from 60.0% to 100.0%; Poppler stayed at 25.0%; ImageMagick remained unsolved.
In FFmpeg, the ddmin reducer, which deletes fixed-size chunks, often violates the media container length and offset invariants and stops reproducing the failure under the same predicate and harness. In contrast, the LLM-generated reducer applies structure-preserving transformations with adaptive chunking to maintain those invariants before driving ddmin, which preserves the crash and still reduces the input.
Table 12. Test case reduction: success rate and compression ratio comparison (Qwen2.5-Plus). SR: Successful Rate, CR: Compress Rate.
Project (n) ddmin-only ReduceFix Pure LLM
SR CR SR CR SR CR
FFmpeg (1) 0.0 N/A 100.0 35.4 0.0 N/A
ImageMagick (1) 100.0 4.6 100.0 5.0 0.0 N/A
MuPDF (5) 100.0 88.1 100.0 87.4 0.0 N/A
PHP-src (1) 100.0 85.6 100.0 97.7 0.0 N/A
Poppler (4) 50.0 22.8 75.0 25.4 0.0 N/A
Overall (12) 75.0 51.8 91.7 56.4 0.0 N/A
Table 13. Pass@K of ReduceFix (Qwen2.5-Plus) on OSS-Fuzz.
Project (n) Baseline Origin Test Reduced Test
@1 @5 @10 @1 @5 @10 @1 @5 @10
FFmpeg (1) 0.0 0.0 0.0 0.0 0.0 0.0 10.0 50.0 100.0
ImageMagick (1) 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
MuPDF (5) 10.0 19.9 20.0 4.0 15.6 20.0 4.0 20.0 40.0
PHP-src (1) 60.0 100.0 100.0 0.0 0.0 0.0 100.0 100.0 100.0
Poppler (4) 25.0 25.0 25.0 22.5 25.0 25.0 25.0 25.0 25.0
Overall (12) 17.5 25.0 25.0 9.2 14.8 16.7 19.2 29.2 41.7
[RQ-5] [Findings]: ReduceFix is effective on repository-level repair scenarios. On OSS-Fuzz, ReduceFix preserves crashes more reliably than ddmin-only while achieving equal or better compression, and those reductions translate into higher pass@k under evaluation. The largest improvements in pass@k occur when ddmin-only fails or compresses poorly, including in FFmpeg and PHP-src. [Insights]: The evaluation demonstrates the effectiveness of ReduceFix in real-world repository-level repair scenarios. By combining semantic-aware LLM reasoning with traditional reduction techniques, ReduceFix reliably preserves failure behavior while producing more concise and interpretable test cases. These reductions substantially improve downstream repair success, confirming that LLM-guided reduction is not merely a preprocessing step but an essential and practical front-end component of automated repair pipelines at the repository level.